OpenStudy (anonymous):

the area of a sphere is 900 π m^2. what is the diameter. i liked it better when it was just circles

OpenStudy (anonymous):

A = 4 pie r^2

OpenStudy (anonymous):

thank you for the equation.

OpenStudy (anonymous):

900 pie = 4 pie r^2 900/4 =r^2 30/2 = r r =15 d = 2r d = 2*15 = 30

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

welcome..)) Plzz read the message which i have send u..

OpenStudy (lgbasallote):

^circle

OpenStudy (anonymous):

ok. im afraid i cant click the link because my mother doesnt want me to -_-